# Hermitian-Singer Functional and Differential Codes

### v1

@@ 1 Introduction
In the projective plane $PG(2,q^{2})$ of order $q^{2}$ equipped with homogeneous coordinates $(X_{1}:X_{2}:X_{0})$ defined over a finite field $\mathbb{F}_{q^{2}}$, the non-singular plane curve of equation $X_{2}^{q}X_{0}+X_{2}X_{0}^{q}-X_{1}^{q+1}=0$ is a canonical form of the Hermitian curve ${\mathcal{H}}(q)$. It is well known that a Hermitian curve is a maximal curve over $\mathbb{F}_{q^{2}}$ with $q^{3}+1$ points in $PG(2,q^{2})$. Take a divisor $\mathtt{G}$ of ${\mathcal{H}}(q)$ with support $G$ where $G$ is a subset of points of ${\mathcal{H}}(q)$, and define $D$ to be the set of all points of ${\mathcal{H}}(q)$ other than those in $G$. Fix an ordering $(P_{1},\ldots,P_{N})$ of the points in $D$ where $N=(q^{3}+1)-|G|$, and define the divisor $\mathtt{D}$ to be the sum $P_{1}+\ldots+P_{N}$. Let $\mathcal{L}(\mathtt{G})$ be the Riemann-Roch space associated with $\mathtt{G}$. An algebraic geometric (shortly AG) code on ${\mathcal{H}}(q)$ arises by evaluating the functions in $\mathcal{L}(\mathtt{G})$ on $D$. Such a code $C_{\mathcal{L}}(\mathtt{D},\mathtt{G})$ is the Hermitian functional code associated to $\mathtt{G}$. The dual of the functional code $C_{\mathcal{L}}(\mathtt{D},\mathtt{G})$ is the differential code $C_{\Omega}(\mathtt{D},\mathtt{G})$, called Hermitian differential code.
The functional codes and their duals on the Hermitian curves may have good performances provided that $\mathtt{G}$ is taken appropriately. One of the best known code of this type is the $1$-point Hermitian functional code where $G$ consists of a single point, i.e. $\mathtt{G}=\lambda Q$ where $\lambda$ is a positive integer, and $Q$ is a point $Q\in{\mathcal{H}}(q)$; see [1] and the references therein. Generalizations of the $1$-point codes have been made so far in two directions. The first consists in replacing $Q$ with a higher degree place of the function field of ${\mathcal{H}}(q)$, which yields differential codes with even better parameters obtained for some values of $\lambda$ such that $q^{3}-q^{2}\leq\lambda\leq q^{3}$; see [3, 15, 16, 19]. The second direction is motivated by the fact that the subgroup $\Gamma$ of the automorphism group $PGU(3,q)$ of ${\mathcal{H}}(q)$ which preserves $G$ is also an automorphism group of $C_{\mathcal{L}}(\mathtt{D},\mathtt{G})$ whenever each point of $G$ appears with the same weight in $\mathtt{G}$.
@@ 2.1 Plane curves and their function fields
In the special case where $\mathbb{L}=\mathbb{F}_{q^{2}}$, $\mathcal{F}$ is the Hermitian curve of $PG(2,q^{2})$, $G=\{Q_{1},\ldots,Q_{q^{2}-q+1}\}$ is a $\Gamma$-orbit of a Singer subgroup of $PGU(3,q)$, and $\mathtt{G}=\lambda(Q_{1}+\ldots+Q_{q^{2}-q+1})$ for a positive integer $\lambda$. Then $\mathfrak{g}=\frac{1}{2}(q^{2}-q)$ and $\deg(\mathtt{G})=\lambda(q^{2}-q+1)$. Since $|G|>2\mathfrak{g}-2$, the Riemann-Roch theorem yields in this case,
@@ 3 Families of Hermitian curves
Let $\delta=\gamma^{q^{2}-q+1}$. Then $1+t\delta^{q+1}+t^{q^{2}+1}\delta=0$ with $\delta^{q^{2}+q+1}=1.$ The equation $tX^{q+1}+t^{q^{2}+1}X+1=0$ has as many as $q+1$ pairwise distinct roots $\varepsilon$ in an algebraic closure of $\mathbb{F}_{q^{2}}$. As $t\varepsilon^{q+1}+t^{q^{2}+1}\varepsilon+1=0$ yields $t^{q}\varepsilon^{q^{2}+q}+t^{q^{3}+q}\varepsilon^{q}+1=0$, we have $t^{q}\varepsilon^{q^{2}+q+1}+t^{q+1}\varepsilon^{q+1}+\varepsilon=0$ whence $t^{q}(\varepsilon^{q^{2}+q+1}+t\varepsilon^{q+1}+t^{q^{2}+1}\varepsilon)=0$ follows. Hence $\varepsilon^{q^{2}+q+1}=1$. Therefore, the elements $\gamma$ satisfying (2) are determined in the following way. First, the $q+1$ solutions $\varepsilon$ of Equation $tX^{q+1}+t^{q^{2}+1}X+1=0$ are computed, all belong to $\mathbb{F}_{q^{3}}$. After that for each such $\varepsilon$ the $q^{2}-q+1$ solutions of Equation $X^{q^{2}-q+1}=\varepsilon$ are computed. All together, $q^{3}+1$ solutions $\gamma$ are obtained. Therefore, $H_{t}$ has exactly $q^{3}+1$ points in $\Pi$. Moreover, $H_{t}$ is defined over $\Pi$ since the Frobenius map $\Phi_{\Pi}$ preserves $H_{t}$. In fact, as $t^{q^{4}+q^{2}}=t^{-1}$,
@@ 4.1 Hermitian Singer functional code
From now on, $\mathbb{L}$ stands for the finite subfield of $\mathbb{K}$ of order $q^{2}$. Fix a $(q^{2}+q+1)$-root of unity $\tau$ in $\mathbb{F}_{q^{6}}$, and let $H_{\tau}$ be a Hermitian curve of $\Pi$ of homogeneous equation $\tau X_{1}X_{2}^{q}+\tau^{q^{2}+1}X_{2}X_{0}^{q}+X_{0}X_{1}^{q}=0$. For $G$, take the orbit of a point $P\in H_{\tau}\cap\Pi$ under the action of a Singer subgroup $\Gamma$ of $PGU(3,q)$. Let $C_{\mathcal{L}}(\mathtt{D},\mathtt{G})$ be the functional code on $H_{\tau}$ where $\mathtt{G}=P_{1}+\ldots P_{N}$ with $G=\{P_{1},\ldots P_{N}\},N=q^{2}-q+1$, and $\mathtt{D}=Q_{1}+\ldots+Q_{n}$ with $D=\{Q_{1},\ldots,Q_{n}\},n=q^{3}+1-(q^{2}-q+1)$ such that $G\dot{\cup}D$ is a partition of the set of all points of $H_{\tau}$ on $\Pi$. We call $C_{\mathcal{L}}(\mathtt{D},\mathtt{G})$ the Hermitian Singer functional code. For every positive integer $\lambda\geq 2$, a generalization is obtained whenever $\mathtt{G}$ is replaced by $\lambda\mathtt{G}$.
@@
The linear system of such polynomials $G(X,Y)$ has dimension $\frac{1}{2}(q^{2}-q)$. Therefore, the functions in (3) form a subspace $\Lambda$ of $\mathcal{L}(\mathtt{G})$ with $\dim(\Lambda)=\frac{1}{2}(q^{2}-q)$. Now, take point $Q\in D$. From Proposition 3.4, there is a Hermitian curve $H_{u}$ through $Q$. Since $\Pi\cap(H_{\tau}\cap H_{t})=G$ and $Q\not\in G$, $H_{u}$ is neither $H_{\tau}$, nor $H_{t}$. Let $H(X,Y)=uXY^{q}+u^{q^{2}+1}Y+X^{q}=0$ be an affine equation of $H_{u}$. Then $H(X,Y)/F(X,Y)\in\mathcal{L}(\mathtt{G})$ and $H(X,Y)/F(X,Y)\not\in\Lambda$. Thus $\Lambda$ together with $H(X,Y)/F(X,Y)$ and a non-zero constant function, determine a basis of $\mathcal{L}(\mathtt{G})$. Since $H(X,Y)/F(X,Y)$ has exactly $q^{2}-q+1$ zeros in $D$, the weight of $H(X,Y)/F(X,Y)$ in $C_{\mathcal{L}}(\mathtt{D},\mathtt{G})$ equals $q^{3}+1-2(q^{2}-q+1)$. On the other hand, the designed minimum distance of $C_{\mathcal{L}}(\mathtt{D},\mathtt{G})$ is $n-\deg(\mathtt{G})=q^{3}+1-2(q^{2}-q+1)$. Therefore, the following result is obtained.
@@
The subspace $\Lambda$ in the above discussion is a subcode of $C_{\mathcal{L}}(\mathtt{D,\mathtt{G}})$ of length $q(q^{2}-q+1)$, dimension $\frac{1}{2}(q^{2}-q)$ and designed minimum distance $q^{3}-2(q^{2}-q-1)$. We show that the real minimum distance equals $q^{3}-2(q^{2}-q)+1$. The zeros of the functions in $\Lambda$ are contained in the intersection of $H_{t}$ and a curve of equation $G(X,Y)$ of degree at most q2q-2. Since $H_{t}$ is absolutely irreducible, the Bézout’s theorem yields that the size of this intersection at most (q+1)(q2)(q+1)(q-2). Actually, this bound is attained for some functions in $\Lambda$. For instance, as $\frac{1}{2}(q-1)^{2}-1\geq q-2$, Proposition 3.9 gives a method to find such a function by taking q2q-2 chords through a suitable chosen point $U\in\Pi\setminus H_{t}$, and define $G(X,Y)$ as the product of the linear polynomials representing those q2q-2 lines. Therefore, the following result holds.
@@ 4.2 Hermitian Singer differential code
From [21, Theorem 12.17], there is a canonical divisor $\mathtt{W}$ such that
@@
where ${\mathtt{U}}=U_{1}+U_{2}+U_{3}$ and $U_{1},U_{2},U_{3}$ are the common points of ${\mathcal{H}}(q)$ and $F$ in the cubic extension $PG(2,q^{6})$ of $PG(2,q^{2})$. As we have already observed in section 4.1, the intersection multiplicity $I(U_{i},{\mathcal{H}}(q)\cap F)$ equals $q$ for $i=1,2,3$, and if $\ell_{1}\ell_{2}\ell_{3}$ is the cubic curve which is the product of the sides $\ell_{i}$ of the triangle $U_{1}U_{2}U_{3}$, then $I(U_{i},{\mathcal{H}}(q)\cap\ell_{1}\ell_{2}\ell_{3})=q+1$ for $i=1,2,3$. From Lemma 3.3,
@@
In the above construction of $f$, if the curve $\mathcal{C}$ happens to have at least $\frac{1}{2}(q-2)(q+1)+\kappa$ common points with $H(q)$ in $PG(2,q^{2})$, then the bound on $d$ in Theorem 4.4 is improved by $\kappa$ so that $3\leq d\leq\frac{1}{2}(q^{2}-q)+2-\kappa$. Unfortunately, finding such curves $\mathcal{C}$ appears to be a rather difficult problem, and no results about it are available yet in the literature.

### v2

@@ Abstract
Keywords: Hermitian Curve; Singer Cycle; AG-Code
MSC: 14H55; 11T71; 11G20; 94B27.
@@ 1 Introduction
In the projective plane $PG(2,q^{2})$ of order $q^{2}$ equipped with homogeneous coordinates $(X_{1}:X_{2}:X_{0})$ defined over a finite field $\mathbb{F}_{q^{2}}$, the non-singular plane curve of equation $X_{2}^{q}X_{0}+X_{2}X_{0}^{q}-X_{1}^{q+1}=0$ is a canonical form of the Hermitian curve ${\mathcal{H}}(q)$. It is well known that a Hermitian curve is a maximal curve over $\mathbb{F}_{q^{2}}$ with $q^{3}+1$ points in $PG(2,q^{2})$, see [25] for more information on the Hermitian curve. Take a divisor $\mathtt{G}$ of ${\mathcal{H}}(q)$ with support $G$ where $G$ is a subset of points of ${\mathcal{H}}(q)$, and define $D$ to be the set of all points of ${\mathcal{H}}(q)$ other than those in $G$. Fix an ordering $(P_{1},\ldots,P_{N})$ of the points in $D$ where $N=(q^{3}+1)-|G|$, and define the divisor $\mathtt{D}$ to be the sum $P_{1}+\ldots+P_{N}$. Let $\mathcal{L}(\mathtt{G})$ be the Riemann-Roch space associated with $\mathtt{G}$. An algebraic geometric (shortly AG) code on ${\mathcal{H}}(q)$ arises by evaluating the functions in $\mathcal{L}(\mathtt{G})$ on $D$. Such a code $C_{\mathcal{L}}(\mathtt{D},\mathtt{G})$ is the Hermitian functional code associated to $\mathtt{G}$. The dual of the functional code $C_{\mathcal{L}}(\mathtt{D},\mathtt{G})$ is the differential code $C_{\Omega}(\mathtt{D},\mathtt{G})$, called Hermitian differential code.
The functional codes and their duals on the Hermitian curves may have good performances provided that $\mathtt{G}$ is taken appropriately. One of the best known code of this type is the $1$-point Hermitian functional code where $G$ consists of a single point, i.e. $\mathtt{G}=\lambda Q$ where $\lambda$ is a positive integer, and $Q$ is a point $Q\in{\mathcal{H}}(q)$; see [1] and the references therein. Generalizations of the $1$-point codes have been made so far in two directions. The first consists in replacing $Q$ with a higher degree place of the function field of ${\mathcal{H}}(q)$, which yields differential codes with even better parameters obtained for some values of $\lambda$ such that $q^{3}-q^{2}\leq\lambda\leq q^{3}$; see [3, 18, 19, 22]. The second direction is motivated by the fact that the subgroup $\Gamma$ of the automorphism group $PGU(3,q)$ of ${\mathcal{H}}(q)$ which preserves $G$ is also an automorphism group of $C_{\mathcal{L}}(\mathtt{D},\mathtt{G})$ whenever each point of $G$ appears with the same weight in $\mathtt{G}$. In such a context, one may also think of replacing the Hermitian curve by some other curves with many points in $PG(2,q)$, for instance by those investigated in [14, 15, 16].
@@
It seems worth investigating whether the geometry developed in this paper might be adequate to compute the minimum distances of the multiple-Hermitian Singer functional codes $C_{\mathcal{L}}(\mathtt{D},\lambda\mathtt{G})$ where $\lambda\geq 2$. This would in turn provide the minimum distances of their dual codes which are multiple-Hermitian Singer differential codes $C_{\Omega}(\mathtt{D},\mu\mathtt{G})$ for $\mu\geq 1$. Also, the expansion of the project including functional and differential codes where $G$ is the union of a few $\Gamma$-orbits may produce some more codes with large automorphism groups.
@@ 2.1 Plane curves and their function fields
In the special case where $\mathbb{L}=\mathbb{F}_{q^{2}}$, $\mathcal{F}$ is the Hermitian curve of $PG(2,q^{2})$, $G=\{Q_{1},\ldots,Q_{q^{2}-q+1}\}$ is a $\Gamma$-orbit of a Singer subgroup of $PGU(3,q)$, and $\mathtt{G}=\lambda(Q_{1}+\ldots+Q_{q^{2}-q+1})$ for a positive integer $\lambda$. Then $\mathfrak{g}=\frac{1}{2}(q^{2}-q)$ and $\deg(\mathtt{G})=\lambda(q^{2}-q+1)$. Since $|G|>2\mathfrak{g}-2$, the Riemann-Roch theorem yields in this case
@@ 3 Families of Hermitian curves
Let $\delta=\gamma^{q^{2}-q+1}$. Then $1+t\delta^{q+1}+t^{q^{2}+1}\delta=0$ with $\delta^{q^{2}+q+1}=1.$ The equation $tX^{q+1}+t^{q^{2}+1}X+1=0$ has as many as $q+1$ pairwise distinct roots $\varepsilon$ in an algebraic closure of $\mathbb{F}_{q^{2}}$. As $t\varepsilon^{q+1}+t^{q^{2}+1}\varepsilon+1=0$ yields $t^{q}\varepsilon^{q^{2}+q}+t^{q^{3}+q}\varepsilon^{q}+1=0$, we have $t^{q}\varepsilon^{q^{2}+q+1}+t^{q+1}\varepsilon^{q+1}+\varepsilon=0$ whence $t^{q}(\varepsilon^{q^{2}+q+1}+t\varepsilon^{q+1}+t^{q^{2}+1}\varepsilon)=0$ follows. Hence $\varepsilon^{q^{2}+q+1}=1$. Therefore, the elements $\gamma$ satisfying (2) are determined in the following way. First, the $q+1$ solutions $\varepsilon$ of Equation $tX^{q+1}+t^{q^{2}+1}X+1=0$ are computed, all belong to $\mathbb{F}_{q^{3}}$. After that for each such $\varepsilon$ the $q^{2}-q+1$ solutions of Equation $X^{q^{2}-q+1}=\varepsilon$ are computed. All together, $q^{3}+1$ solutions $\gamma$ are obtained. Therefore, $H_{t}$ has exactly $q^{3}+1$ points in $\Pi$. Moreover, $H_{t}$ is defined over $\Pi$ since the Frobenius map $\Phi_{\Pi}$ preserves $H_{t}$. In fact, since $\Phi_{\Pi}(x_{1})=x_{0}^{q},\Phi_{\Pi}(x_{2})=x_{1}^{q},\Phi_{\Pi}(x_{0})=x_{2}^{q}$ and $t^{q^{4}+q^{2}}=t^{-1}$, we have
@@ 4.1 Hermitian Singer functional code
From now on, $\mathbb{L}$ stands for the finite subfield of $\mathbb{K}$ of order $q^{2}$. Fix a $(q^{2}+q+1)$-th root of unity $\tau$ in $\mathbb{F}_{q^{6}}$, and let $H_{\tau}$ be a Hermitian curve of $\Pi$ of homogeneous equation $\tau X_{1}X_{2}^{q}+\tau^{q^{2}+1}X_{2}X_{0}^{q}+X_{0}X_{1}^{q}=0$. For $G$, take the orbit of a point $P\in H_{\tau}\cap\Pi$ under the action of a Singer subgroup $\Gamma$ of $PGU(3,q)$. Let $C_{\mathcal{L}}(\mathtt{D},\mathtt{G})$ be the functional code on $H_{\tau}$ where $\mathtt{G}=P_{1}+\ldots P_{N}$ with $G=\{P_{1},\ldots P_{N}\},N=q^{2}-q+1$, and $\mathtt{D}=Q_{1}+\ldots+Q_{n}$ with $D=\{Q_{1},\ldots,Q_{n}\},n=q^{3}+1-(q^{2}-q+1)$ such that $G\dot{\cup}D$ is a partition of the set of all points of $H_{\tau}$ on $\Pi$. We call $C_{\mathcal{L}}(\mathtt{D},\mathtt{G})$ the Hermitian Singer functional code. For every positive integer $\lambda\geq 2$, a generalization is obtained whenever $\mathtt{G}$ is replaced by $\lambda\mathtt{G}$.
@@
The linear system of such polynomials $G(X,Y)$ has dimension $\frac{1}{2}(q^{2}-q)$. Therefore, the functions in (3) form a subspace $\Lambda$ of $\mathcal{L}(\mathtt{G})$ with $\dim(\Lambda)=\frac{1}{2}(q^{2}-q)$. Now, take point $Q\in D$. By Proposition 3.4, there is a Hermitian curve $H_{u}$ through $Q$. Since $\Pi\cap(H_{\tau}\cap H_{t})=G$ and $Q\not\in G$, $H_{u}$ is neither $H_{\tau}$, nor $H_{t}$. Let $H(X,Y)=uXY^{q}+u^{q^{2}+1}Y+X^{q}=0$ be an affine equation of $H_{u}$. Then $H(X,Y)/F(X,Y)\in\mathcal{L}(\mathtt{G})$ and $H(X,Y)/F(X,Y)\not\in\Lambda$. Thus $\Lambda$ together with $H(X,Y)/F(X,Y)$ and a non-zero constant function, determine a basis of $\mathcal{L}(\mathtt{G})$. Since $H(X,Y)/F(X,Y)$ has exactly $q^{2}-q+1$ zeros in $D$, the weight of $H(X,Y)/F(X,Y)$ in $C_{\mathcal{L}}(\mathtt{D},\mathtt{G})$ equals $q^{3}+1-2(q^{2}-q+1)$. On the other hand, the designed minimum distance of $C_{\mathcal{L}}(\mathtt{D},\mathtt{G})$ is $n-\deg(\mathtt{G})=q^{3}+1-2(q^{2}-q+1)$. Therefore, the following result is obtained.
@@
The subspace $\Lambda$ in the above discussion is a subcode of $C_{\mathcal{L}}(\mathtt{D,\mathtt{G}})$ of length $q(q^{2}-q+1)$, dimension $\frac{1}{2}(q^{2}-q)$ and designed minimum distance $q^{3}-2(q^{2}-q-1)$. We show that the real minimum distance equals $q^{3}-2(q^{2}-q)+1$. The zeros of the functions in $\Lambda$ are contained in the intersection of $H_{t}$ and a curve of equation $G(X,Y)$ of degree at most q2q-2. Since $H_{t}$ is absolutely irreducible, the Bézout theorem yields that the size of this intersection at most (q+1)(q2)(q+1)(q-2). Actually, this bound is attained for some functions in $\Lambda$. For instance, as $\frac{1}{2}(q-1)^{2}-1\geq q-2$, Proposition 3.9 gives a method to find such a function by taking q2q-2 chords through a suitable chosen point $U\in\Pi\setminus H_{t}$, and define $G(X,Y)$ as the product of the linear polynomials representing those q2q-2 lines. Therefore, the following result holds.
@@ 4.2 Hermitian Singer differential code
From [24, Theorem 12.17], there is a canonical divisor $\mathtt{W}$ such that
@@
where ${\mathtt{U}}=U_{1}+U_{2}+U_{3}$ and $U_{1},U_{2},U_{3}$ are the common points of ${\mathcal{H}}(q)$ and $F$ in the cubic extension $PG(2,q^{6})$ of $PG(2,q^{2})$. As we have already observed in section 4.1, the intersection multiplicity $I(U_{i},{\mathcal{H}}(q)\cap F)$ equals $q$ for $i=1,2,3$, and if $\ell_{1}\ell_{2}\ell_{3}$ is the cubic curve which is the product of the sides $\ell_{i}$ of the triangle $U_{1}U_{2}U_{3}$, then $I(U_{i},{\mathcal{H}}(q)\cap\ell_{1}\ell_{2}\ell_{3})=q+1$ for $i=1,2,3$. By Lemma 3.3,
@@
In the above construction of $f$, if the curve $\mathcal{C}$ happens to have at least $\frac{1}{2}(q-2)(q+1)+\kappa$ common points with $H(q)$ in $PG(2,q^{2})$, then the bound on $d$ in Theorem 4.4 is improved by $\kappa$ so that $3\leq d\leq\frac{1}{2}(q^{2}-q)+2-\kappa$. Unfortunately, finding such curves $\mathcal{C}$ appears to be a rather difficult problem, and no results about it are available in the literature yet.